Use the calculus to find the approximate change in the volume of a spherical soap bubble when its radius increases from 3 cm to 3.01 cm. What is the approximate value of the volume when the radius is 3.01 cm?


1
Expert's answer
2021-03-24T15:02:41-0400

"V={4 \over 3}\pi r^3"


When the radius was 3cm the volume was "{4 \over 3}×3.14 ×3^3=113.04cm^3"


When the radius is now 3.01

dr= 0.01cm


"dV={4 }\pi r^2dr"

"dV=4\u00d73.14\u00d73^2\u00d70.01=1.1304cm^3"


The volume when the radius is now "3.01cm=(113.04 + 1.1304)cm^3= 114.17cm^3"